OpenStudy (anonymous):

The Okarche water department has a monthly service charge of $14.60 and a volume charge of $1.13 for every 100 cubic feet of water used. Which of the following equations can be used to determine the Savage family's monthly water bill? (Let x represent the amount of water used, measured as indicated above, and y represent the monthly cost.)

OpenStudy (anonymous):

@anusha.p @Callisto ?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

could you help with the question

OpenStudy (callisto):

Total amount = service charge + volume charge total amount =y service charge = 14.60 volume charge = (x/100 )(1.13) Therefore, y= 14.60 + (x/100 )(1.13) = 14.60 + (1.13x/100) Can you further simplify it?

OpenStudy (anonymous):

@callisto,if the volume usage is less than 100 cubic feet.he won't be charged 1.13 dollars i thinkkk.

OpenStudy (anonymous):

y = 1.13x + 14.60 y = 0.0113x + 14.60 y = 15.73x y = 1.13x - 14.60 so which would it be

OpenStudy (callisto):

1.13 per every 100 cubic meter so, 1.13/100 per cubic meter, that is 0.0113 per cubic meter He uses x cubic meter so volume charge = $0.0113x @anusha.p I haven't said he will be charged 1.13 dollars if the volume usage is less than 100 cubic feet :|

OpenStudy (callisto):

@selenamalter you cannot add the 2 terms :|

OpenStudy (anonymous):

so its B

OpenStudy (anonymous):

of course @Callisto .i just got confusedd,if it is not charged when the usage is less than 100 ft^3.

OpenStudy (callisto):

@selenamalter What do you mean? BTW, is the vol. of water measured in 100 cub. meter?

OpenStudy (anonymous):

i sent the answer options

OpenStudy (callisto):

@anusha.p mathematically, that is the equation... I don't know what the real case is, as I'm not the one who responsible for paying the bills..

OpenStudy (anonymous):

K.LOL

OpenStudy (callisto):

@selenamalter you haven't mentioned there are choices. but it must not be -14.60 ... it should be + 14.60 :|